OpenStudy (campayne):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) {2x - y = 2 3x - 2y = 11

OpenStudy (mathmale):

If you must use substitution, choose whichever of the 2 given equations can be more easily solved for either x or y. Then, substitute your result into the other equation and solve for the remaining variable. Show your work.

OpenStudy (will.h):

still needs help mate?

OpenStudy (campayne):

yes i do

OpenStudy (will.h):

2x - y = 2 3x - 2y = 11 That's the system of equations you have you have to use substitution method correct? don't answer that we know already so anyway one of those equations MUST have only one variable on one side the easiest is to isolate y for the 1st equation let's see 2x - y = 2 isolate y -y = 2 - 2x y = -2 + 2x so now mate we have y = -2 + 2x 3x - 2y = 11 now we will SUBSTITUTE y value in the 2nd equation 3x -2(-2 + 2x) = 11 Now solve this for x

OpenStudy (mathmale):

Even if your first efforts are wrong, you'd learn more by trying to solve this problem than by doing nothing. Hint: 2x - y = 2 is easier to solve for y than is the other equation. So, solve 2x - y = 2 for y. You'll get a 2-term expression in x. Substitute this expression for y in the other equation. Will.H has explained this well, already. Please share your work now.

OpenStudy (will.h):

let me know if you need additional help

OpenStudy (mathmale):

Cam: When you have others helping you, it's most polite to respond right away.

OpenStudy (campayne):

right right sorry im still working this out the math

OpenStudy (campayne):

okay so so far ive got - 14 and -16

OpenStudy (will.h):

see @mathmale is a really wonderful professor

OpenStudy (will.h):

he'll take it from here have to help others

OpenStudy (mathmale):

Cam: Easier for the rest of us to help you meaningfully if you show how you got your results. Thanks, Will, for the lovely compliment.

OpenStudy (will.h):

my pleasure

OpenStudy (mathmale):

If you haven't already, solve 2x - y = 2 for y. Hint: This might be easier if you change the signs of all 3 terms.

OpenStudy (campayne):

well wouldnt that be y = -2 + 2x

OpenStudy (mathmale):

Yes, very good. Now, look at the second equation:

OpenStudy (mathmale):

3x - 2y = 11 Enclose the "y" in parentheses: 3x - 2( y )= 11 Now throw out "y" and replace it with your "-2+2x."

OpenStudy (campayne):

3x - 2 (y = -2 + 2x) = 11

OpenStudy (mathmale):

Yes, but OMIT the "y=" part completely. Your result?

OpenStudy (mathmale):

y is GONE. So is the extra "=" sign.

OpenStudy (campayne):

3x - 2 (-2 + 2x) = 11

OpenStudy (mathmale):

You're fast, and that's correct. Please multiply out 12(-2+2x) now. What is the simplified form of your 3x - 2 (-2 + 2x) = 11 ?

OpenStudy (mathmale):

I'm sorry, that's "multiply out 2(-2+2x)."

OpenStudy (mathmale):

You should get something like this: 3x + ? - ? = 11.

OpenStudy (campayne):

im lost now.. multiply out?

OpenStudy (mathmale):

-2(-2+2x) implies MULTIPLICATION.

OpenStudy (mathmale):

What's -2 times -2? What's -2 times 2x? What's -2 times (-2+2x)?

OpenStudy (campayne):

4 + 4x

OpenStudy (mathmale):

Wouldn't that be 4-4x?

OpenStudy (campayne):

oh yeah true

OpenStudy (mathmale):

Given 3x - 2( y )= 11, replace -2(y) with your 4-4x.

OpenStudy (campayne):

3x (4 - 4x) = 11

OpenStudy (mathmale):

Missing sign. 3x +4 -4x = 11 Combine 3x -4x. Move that '4' to the right side (use subtraction to do that).

OpenStudy (campayne):

x - 4x = 11

OpenStudy (mathmale):

Have to subtract 4 from BOTH sides: 3x-4x = 11-4 Simplify this. Find x.

OpenStudy (campayne):

so.. 3x - 11?

OpenStudy (mathmale):

3x-4x is -x, and 11-4 is 7. You should end up with -x=7. Solve this for x.

OpenStudy (campayne):

well -7

OpenStudy (mathmale):

x=-7.

OpenStudy (campayne):

yes..

OpenStudy (mathmale):

Next, we have to find the value of y. The 2nd equation is 3x - 2y = 11. Replace "x" here with (-7). Be sure to include those parentheses. Find y

OpenStudy (campayne):

-21 - 2y = 11

OpenStudy (mathmale):

Very good. Must now solve this for y. To do this, please ADD 21 to both sides of the equation.

OpenStudy (campayne):

so -2y = 31? or -21 + 19 = 11

OpenStudy (mathmale):

-21 - 2y = 11 +21 =+21 ----------------- 0 -2y = 32 find y

OpenStudy (campayne):

-16

OpenStudy (mathmale):

Correct! Sorry this discussion seems so long. I am going to go thru this problem really quickly: {2x - y = 2 3x - 2y = 11 Solving the first eq'n for y, -y=2-2x, or y=2x-2. Subbing this last result into the 2nd equation, 3x-2(2x-2)=11

OpenStudy (mathmale):

3x-4x+4=11 -x =7, or x=-7

OpenStudy (mathmale):

Just wanted to show you how fast this goes once we've gotten the hang of it.

OpenStudy (mathmale):

{2x - y = 2 3x - 2y = 11 2nd equation is 3x-2y=11. Find y. Subbing x=-7, 3(-7) - 2y = 11 -21 -2y = 11 y=-16 Solution is ( x , y) Please write in the values of x and y that we've just found.

OpenStudy (mathmale):

solution is (

OpenStudy (campayne):

(-7, -16)

OpenStudy (mathmale):

Cool. Done. Questions?

OpenStudy (campayne):

no for now no imma need to look over what we did and thank you :)

OpenStudy (mathmale):

We were using the .... method. Can you remember?

OpenStudy (campayne):

substitution right

OpenStudy (mathmale):

yes, right! good! we subbed y=-4+4x. Good for you! Til later, then, bye.